specify all x & y- intercepts of this equation:
y=(x-1)^2+4

Answer by stanbon(75887)

specify all x & y- intercepts of this equation:
y=(x-1)^2+4
--------------------------
y-intercept:
Let x = 0
y = (0-1)^2+4 = 1+4 = 5
----------------------------
x-intercept:
Let y = 0
(x-1)^2+4 = 0
(x-1)^2 = -4
x-1 = 2i or x-1 = -2i
x = 1+2i or x = 1-2i
These are zeroes of the function,
but they are not x-intercepts
because they are not Real Numbers.
============================
Cheers,
Stan H.
============================